### v1

# The Geometry of Queries: Query-Based Innovations in Retrieval-Augmented Generation

Abstract
Digital health chatbots powered by Large Language Models (LLMs) have the potential to significantly improve personal health management for chronic conditions by providing accessible and on-demand health coaching and question-answering. However, these chatbots risk providing unverified and inaccurate information because LLMs generate responses based on patterns learned from diverse internet data. Retrieval Augmented Generation (RAG) can help mitigate hallucinations and inaccuracies in LLM responses by grounding it on reliable content. However, efficiently and accurately retrieving most relevant set of content for real-time user questions remains a challenge.
In this work, we introduce Query-Based Retrieval Augmented Generation (QB-RAG), a novel approach that pre-computes a database of potential queries from a content base using LLMs. For an incoming patient question, QB-RAG efficiently matches it against this pre-generated query database using vector search, improving alignment between user questions and the content. We establish a theoretical foundation for QB-RAG and provide a comparative analysis of existing retrieval enhancement techniques for RAG systems. Finally, our empirical evaluation demonstrates that QB-RAG significantly improves the accuracy of healthcare question answering, paving the way for robust and trustworthy LLM applications in digital health.
Keywords: RAG, LLM, Healthcare, Digital Health.
11footnotetext: $\dagger$ Equal contribution.
1 Introduction
Large Language Models (LLMs), such as OpenAI’s GPT (OpenAI, 2024), Meta’s Llama (Touvron et al., 2023) and Google’s Palm (Anil et al., 2023) and Gemini (Team et al., 2023) suite, have demonstrated remarkable capabilities across a diverse set of natural language understanding and generation tasks. This has led to several novel applications of language models across different domains. In healthcare, LLMs hold immense promise for developing conversational AI systems that can answer patient questions, offer personalized health advice, and potentially improve access to care, particularly for underserved populations (Clusmann et al., 2023; Peng et al., 2023; Thirunavukarasu et al., 2023; Alowais et al., 2023; Nori et al., 2023; Singhal et al., 2022, 2023; Tu et al., 2024).
However, applying LLMs in healthcare presents significant challenges in ensuring the accuracy, reliability, safety, and adherence to the latest healthcare practices in the information they provide. The probabilistic nature of LLMs, combined with limitations and potential biases in their training data, can lead to the generation of fabricated or nonsensical information, known as hallucinations. This poses significant risks in a healthcare context where accurate information is paramount. Additionally, the knowledge embedded within an LLM is limited to the data it was trained on and may not reflect the most current medical guidelines and best practices. To address these challenges, further fine-tuning, instruction tuning (Wei et al., 2022) and additional reinforcement learning (Ouyang et al., 2022) approaches can be considered. Common approaches involve training the models on specific datasets containing the most up-to-date information and/or preferred responses. However, fine-tuning has its limitations: datasets can be difficult and expensive to acquire, especially in healthcare and the computation cost may be prohibitive.
Retrieval Augmented Generation (RAG) (Lewis et al., 2020) offers a promising alternative by grounding LLM responses in a curated and reliable knowledge base of vetted information. Unlike relying solely on the LLM’s internal knowledge, RAG systems retrieve relevant information for a specific user query from this external knowledge base to inform the LLM’s answer generation process. This grounding helps to mitigate hallucinations, promotes factual accuracy, and allows developers to incorporate the latest knowledge. Applications using RAG systems have been shown to be quite compelling compare to traditional LLM fine tuning (Gupta et al., 2024; Ovadia et al., 2024). However, the effectiveness of RAG hinges critically on its ability to accurately retrieve the most pertinent information from the knowledge base, a task complicated by the inherent semantic gap between natural language user queries and the way information is structured and stored within a knowledge base (Ma et al., 2023). This “retrieval challenge” is a significant bottleneck for building effective RAG systems.
Several approaches have attempted to overcome this retrieval challenge by improving the alignment between user queries and the knowledge base. Some methods leverage the generative capabilities of LLMs themselves. For instance, Query2Doc (Wang et al., 2023) and HyDE (Gao et al., 2022) propose generating a hypothetical document that would ideally answer the user’s query. This synthetic document, more semantically aligned with the knowledge base content, is then used as the retrieval key. Similarly, QA-RAG (Kim and Min, 2024) takes a two-pronged approach: first generating a candidate answer using a fine-tuned LLM, and then leveraging both the initial query and this candidate answer to retrieve relevant information from the knowledge base. (Ma et al., 2023) proposes duplicating and/or splitting user queries and then retrieving content similar to either rewritten query. Alternatively, some methods focus on fine-tuning embedding models to create semantically aligned representations of queries and documents (Karpukhin et al., 2020). Despite these efforts, efficiently bridging the semantic gap between user queries and knowledge bases for accurate and reliable information retrieval in RAG systems remains an active area of research.
1.1 Contributions
This paper introduces a novel framework for analyzing and enhancing the retrieval phase of RAG systems. We propose that effective retrieval hinges on aligning user queries with the knowledge base across distinct semantic representations: query, answer, and content. To the best of our knowledge, we are the first to explicitly recognize these distinct “semantic spaces” for improving retrieval. While prior methods have explored alignment based on answer or content representations, our work proposes direct alignment within the query space itself. Existing approaches often implicitly combine these distinct representations, leading to sub-optimal retrieval. In contrast, we systematically disentangle these semantic spaces and demonstrate the advantages of our query based alignment strategy. Specifically, we make the following contributions:
  1. 1.
    QB-RAG: Query-Based Retrieval Augmented Generation: We propose QB-RAG, a novel retrieval method for RAG systems that addresses the misalignment between user queries and content bases by pre-generating a comprehensive set of questions from the knowledge base. Distinct from the methods that rely on generate hypothetical documents or candidate answers, QB-RAG directly maps incoming queries to this offline-generated question set using efficient vector search, enabling efficient, direct query-to-query comparison for enhanced retrieval.
  2. 2.
    Theoretical Foundation for Query-Space Alignment: We develop a theoretical framework that formalizes the retrieval challenge in RAG systems and elucidates the advantages of query-space alignment. Our analysis underscores the inherent semantic gap between user queries and content embeddings, demonstrating how QB-RAG effectively bridges this gap by operating directly within the query space.
  3. 3.
    Survey and Comparative Analysis of Retrieval Strategies: We present a comprehensive survey of prominent retrieval strategies for RAG systems, including naive RAG, QA-RAG, HyDE, and our proposed QB-RAG method. We analyze how each approach addresses the semantic gap between user queries and the knowledge base, comparing their effectiveness through quantitative evaluation on our healthcare dataset. This analysis provides insights into the trade-offs involved in selecting a particular retrieval method for a specific application.
  4. 4.
    Comprehensive Benchmark and Empirical Validation: To evaluate the performance of various retrieval methods, we conduct rigorous experiments on a healthcare question-answering dataset generated from our proprietary healthcare content database. Our evaluation encompasses both retrieval accuracy (using metrics like exact recovery rate and relevance scores) and the quality of the downstream generated answers (assessed using faithfulness, relevancy, and accuracy metrics). This comprehensive benchmark provides valuable insights into the strengths and limitations of each retrieval strategy, showcasing QB-RAG’s superiority on our question-answering dataset. While our evaluation focuses on healthcare question-answering, QB-RAG’s underlying principles and advantages may be broadly applicable to various domains and RAG applications.
The remainder of this paper is structured as follows. Section 2 introduces our healthcare-focused content base and details our methodology for extracting a comprehensive set of questions from these materials. In Section 3, we formalize the RAG objective, highlighting the challenges of retrieval alignment. We then present QB-RAG as a solution, theoretically motivating its effectiveness in bridging the semantic gap between queries and content. Section 4 outlines our benchmark setup, encompassing a range of prominent retrieval methods and evaluation metrics. Finally, Section 5 presents the results of our empirical evaluation, demonstrating QB-RAG’s superiority in retrieval accuracy and downstream answer quality.
2 Application and Dataset
2.1 Description
Digital health programs represent a paradigm shift in managing chronic conditions such as Type 2 diabetes (T2D) and Hypertension (HTN). They can improve health markers associated with chronic conditions by delivering personalized care and support directly to patients (Majithia et al. (2020)). These programs often utilize a variety of digital tools, including mobile apps, messaging, wearable devices, and telehealth services, to facilitate remote monitoring, provide educational content, and offer guidance on lifestyle modifications. The overarching goal is to equip patients with the knowledge and tools necessary to effectively self-manage their condition, enhance their overall quality of life, and potentially reduce the need for frequent traditional healthcare visits.
An integral component of these digital health programs is the messaging platform they incorporate. This platform is not merely a channel for communication; it is an integrated system designed to provide timely, personalized, and interactive support to patients. The effectiveness of these platforms hinges on their ability to engage patients and provide relevant, timely information in response to their queries. However, the increasing demand for immediate healthcare information often strains the capacity of human healthcare providers to deliver timely support through messaging.
To overcome this scalability challenge, digital health platforms are increasingly turning to meticulously curated content repositories designed to provide readily available answers to common patient questions. This study utilizes such a content repository, referred to as “Content Cards,” developed to address frequently asked questions (FAQs) from patients with T2D and HTN participating in our proprietary digital health programs. This dataset, encompassing 630 English-language content cards, covers a comprehensive range of topics pertinent to managing T2D and HTN. This includes general health information, detailed guidance on using platform-specific features and connected devices (e.g., blood glucose monitors), and personalized dietary recommendations. Figure 1 illustrates how these Content Cards are presented within our mobile application, while the excerpt below provides a representative sample of the content:
If there is not enough blood on the test strip, you may not get an accurate blood glucose reading. Some meters won’t even give you a reading at all. So here are some tips to help you get a big enough drop of blood: Rub your hands in warm water to get the blood to your fingertips. Shake your hand to help force blood to your fingertips. Hold your hand down by your side for 30 seconds to help blood run to your fingertips. Set your lancing device to puncture just deep enough to get the size drop you need. (This may take some trial and error.) Take blood from the side of your finger. There are fewer nerves there, so it doesn’t hurt as much. If your lancet isn’t going deep enough, dial the number up higher on the lancing device. Once you’ve lanced, squeeze your finger where it meets your palm and move toward the tip of your finger.
Figure 1: Illustration of Content Cards in Verily’s Onduo mobile application, these cover broad range of topics, from operational health management, nutrition and physical advice, etc.
2.2 Offline Question Generation
The core premise of our work is to improve content retrieval in RAG systems by directly aligning incoming user queries with a pre-computed set of questions derived from our content base. This requires generating a comprehensive set of questions that are answerable by our content, which we achieve through a two-step process. First, we employ a base prompt to generate a preliminary set of questions for each content card. Subsequently, we leverage an LLM-based answerability model (detailed in Section 2.2.2) to filter out irrelevant or nonsensical questions, ensuring the quality and relevance of our question set.
2.2.1 Base prompt
To generate an initial set of questions, we design a prompt (shown in listing 1) with clear instructions for question generation and few-shot examples to illustrate the desired style and format. This prompt incorporates a num_questions parameter to specify the target number of questions per content card. While we set this parameter to 20, the actual number of questions generated by the LLM may vary. Applying this prompt to our content base of 630 cards resulted in over 8,000 potential questions.
1***SYSTEM:***
2You are a Teacher / Professor. Your task is to setup {num_questions} questions for an upcoming quiz/examination. The questions should be both diverse and exhaustive in nature across the document. Restrict the questions to the context information provided.
3
4***INSTRUCTIONS:***
5You are presented with a text authored by healthcare professionals, offering advice and strategies for managing conditions such as diabetes and high blood pressure. Your task is to formulate relevant questions that the text is written to address. Closely follow the example questions for style and structure when formulating your own question for the provided text. Your generated questions should be in first person with details, but only at a high school reading level. Your questions should be answerable from the text, but do not copy the text verbatim. MAKE SURE to generate at least {num_questions} questions. Format the generated questions separated by comma in the following JSON format with "questions" as its key: {{"questions": ["...","...","...","..."]}}
6
7***EXAMPLE OUTPUTS:***
8Sample Text: {example content}
9Generated Questions: {example JSON with list of num_question questions}
10
11--------------------------------------------
12Given the context information and no prior knowledge, generate the relevant questions.
13Text:
14{cc_text}
15Generated Questions:
Listing 1: Base prompt for question generation
2.2.2 Answerability model
After generating questions using an LLM for our content cards, we noticed that our output questions were sometimes not relevant to the content, or were formatted inappropriately. To mitigate this issue and filter out irrelevant questions, we developed an LLM-based answerability model. Specifically, for each content card and generated question pair, we prompt the LLM to assess whether or not the content contains relevant information to answer the question. Leveraging the LLM’s reasoning abilities, we structure the prompt to elicit a step-by-step explanation of the answerability judgment. The LLM is instructed to first identify the key information required to answer the question and then determine if that information is present within the content. The LLM also outputs a “Yes” or “No” label which we use to filter out questions that are deemed irrelevant to the corresponding content card.
Applying this answerability model effectively filters out irrelevant or nonsensical questions generated in the initial phase. This filtering process resulted in a refined set of approximately 4,800 answerable questions derived from our 630 content cards. Each question in this curated set is directly mapped back to its source content card, which serves as the “golden” source for a correct answer.
To validate the effectiveness and reliability of our answerability model, we randomly selected 100 question and content card pairs and evaluated them using both our automated model and three clinical experts. The clinical experts rated the answerability of each pair on a three-point scale: “Content answers completely,” “Content answers partially,” or “Content does not answer.” The results demonstrated strong agreement between our model’s assessments and the clinical expert’s judgments. Specifically, 90% of the pairs received the same answerability rating, with an additional 9% categorized as “partially answerable,” indicating some degree of inherent subjectivity in the evaluation task.
This high level of agreement, coupled with the increasing recognition of LLMs as reliable auto-evaluation tools Lee et al. (2023), supports the validity of our answerability model for filtering irrelevant questions. We provide the prompt used for our answerability model in Listing 2.
1***SYSTEM***
2You are a health coach providing support for members living with diabetes. You have some basic healthcare and nutrition knowledge.
3
4***INSTRUCTIONS***
5Given a pair of user query and a paragraph of content, determine if the content contains relevant information to infer an answer to the query. Think step by step. First provide an explanation, then generate a "Yes" or "No" label. Put the results in a Python dictionary format with keys "Explanation" and "Source relevant".
6
7***EXAMPLES***
8For example, given the following query and content as inputs: {positive and negative examples, with explanations}
9
10--------------------------------------------
11Provide the output for the following query and content:
12Question: {question}
13Content: {content}
Listing 2: Prompt for Answerability Auto-Evaluator
2.2.3 Analysis of Generated Question Similarity
To assess the potential of our question-based retrieval approach, we analyzed the semantic similarity among the questions generated using the prompt in Listing 1. Our goal was to determine if questions derived from the same content card (intra-content similarity) exhibit higher similarity than questions generated from different cards (inter-content similarity).
We calculated pairwise cosine similarity between question embeddings generated using Google’s textembedding-gecko model. As shown in Table 1, mean intra-content similarity (0.871) is indeed higher than mean inter-content similarity (0.827). Similarly, questions are, on average, more similar to their source content card (0.837) than to other content cards (0.762). We note that the relatively high similarity between questions and content from different cards likely stems from the inherent thematic overlap within our dataset, as all content and questions focus on T2D and HTN management.
These results provide initial evidence supporting the premise of our question-based retrieval strategy. While the specific similarity values are influenced by our dataset and embedding model, the consistent trend of higher intra-content similarity suggests that aligning incoming user queries with this question set can improve retrieval accuracy in RAG systems.
Average Cosine Similarity b/w Value
Questions generated from the same content .871
Questions generated from a different content .827
Questions and their associated content .837
Questions and other contents .762
Table 1: Cosine similarities between questions and contents combination
3 Mathematical Formulation
This section presents a mathematical framework for analyzing retrieval within RAG systems, highlighting the limitations of conventional approaches and motivating our proposed QB-RAG method. We formalize the retrieval objective, elucidating the misalignment between user queries and content representations. Subsequently, we introduce QB-RAG and provide a theoretical rationale for its efficacy in addressing this misalignment.
3.1 Notation
Consider the set of $M$ content documents $\mathcal{C}=\{c_{1},\cdots,c_{M}\}$ appropriately chunked for embedding. Similarly, let $\mathcal{Q}=\{q_{1},\cdots,q_{N}\}$ represent the set of $N$ questions generated from the content documents as described in section 2.2. Unless otherwise specified, $c\in\mathcal{C}$ and $q\in\mathcal{Q}$ will refer to their respective embedding representations.
We note that in much of the RAG related literature, the embedder for the query and content base are often the same. As mentioned, this typically leads to misaligned retrieval when comparing query embeddings to content embeddings through cosine similarity. Given cosine similarity as the metric used in our retrieval experiments, we will assume without loss of generality that all embeddings are normalized, i.e. $\forall c\in\mathcal{C},\forall q\in\mathcal{Q}:\>\lVert c\rVert_{2}=\lVert q\rVert_{2}=1$. We also define the distance function as the cosine distance $d(x,y)=1-\frac{x^{t}y}{\lVert x\rVert\lVert y\rVert}$.
The question generation process in section 2.2 is a one to many process, that is from one content we generate multiple questions. To encode this relationship, we denote the matrix $A\in\{0,1\}^{M\times N}$ where $A_{ij}=\mathds{1}[c_{i}\text{ generated }q_{j}]$, where $\mathds{1}$ denotes the indicator function. Per our generation process and the relevance evaluation, it is understood that $A_{ij}=1\Rightarrow q_{j}\text{ can be answered using }c_{i}$. However the converse is not necessarily true as potentially different contents may be able to answer the same question. Thus we define $A^{*}\in\{0,1\}^{M\times N}$ which is the dense unobserved matrix such that $A^{*}_{ij}=\mathds{1}[q_{j}\text{ can be answered using }c_{i}]$. This implies that $A$ is a partial observation of $A^{*}$, which we call the oracle matrix. We point out that following our question generation process described in section 2, we have $\forall c_{j}\in\mathcal{C}:\>\exists q_{i}\in\mathcal{Q}\text{ s.t. }A_{ij}=1$.
For a new user question $q_{0}$, traditional RAG systems retrieve content that maximize some measure of similarity $\arg\max_{c\in\mathcal{C}}{c^{t}q_{0}}$ (or equivalently $\arg\min_{c\in\mathcal{C}}d(c,q_{0})$). More advanced RAG systems rely on a specific functional form $f$, ideally aligned with the objective of downstream generation and retrieves content maximizing $\arg\max_{c\in\mathcal{C}}f({c^{t},q_{0}})$. We acknowledge that in general, even for traditional RAGs, retrieving such a content may be challenging especially for very large content bases. This has led to various key highly efficient retrieval algorithms (e.g. ScaNN111 https://blog.research.google/2020/07/announcing-scann-efficient-vector.html ), and our work takes the above as a solved problem. Further in RAG systems, we usually retrieve a subset $\mathcal{S}\subset\mathcal{C}$ content pieces of size $\lvert\mathcal{S}\rvert=k$ solving $\arg\min_{\mathcal{S}}\sum_{c\in\mathcal{S}}{d(c,q_{0})}$. Oftentimes, different retrieval methods encourage the retrieved set to be diverse to improve the downstream retrieval; these ad-hoc methods are not the focus of our work.
For simplicity in our mathematical formulation, we will assume the retrieval to mean a single piece of content in this section, where RAG retrieves the most similar content to the incoming user query. We use Google’s textembedding-gecko in the current work for generating embeddings. We denote the embedding matrices for content base and questions, respectively, with capital letters $C\in\mathbb{R}^{d\times M}$ and $Q\in\mathbb{R}^{d\times N}$ where $d$ is the dimension of the chosen embedder, in our case 768.
As a general note, we usually denote sets as calligraphic, vectors as lower case, matrices as upper case, and numbers/indices lower or upper case.
3.2 Ideal Retrieval Objective
Ideally during the retrieval phase, for a new user query $q_{0}$, one would evaluate all documents $c\in\mathcal{C}$ by asking the retrieval question “Can the query $q_{0}$ be answered using the content $c$?”. We call this the retrieval task and define the retrieval function $f^{*}(c,q_{0})\in\{0,1\}$. Note this is the exact formulation of the matrix $A^{*}$ for the generated questions, i.e. $A^{*}=(f^{*}(c_{i},q_{j}))_{ij}$.
State-of-the-art LLMs have shown their ability to accurately approximate the retrieval function (which we validate in our experimental section), which means that we can in-principle solve the retrieval task above at scale for any incoming query. However, in most practical application, using LLMs for approximating the retrieval function is both time and cost prohibitive, as it would involves $M$ LLM predictions for an online user. This necessitates exploring alternate schemes.
Rather than pursuing this exact operation for every entry, RAG systems typically approximate the retrieval task with some proxy. Below we list a few of the most common attempts at approximating the retrieval task at scale. The underlying challenge is how should one best approximate the retrieval function using practically scalable operations, for instance, using embedding similarities.
  • The vanilla version of RAG approximates $f^{*}(c,q_{0})$ (and in particular its induced ranking) by $c^{t}q$. We will show below that this approximation, albeit very wide spread, is typically misaligned and can lead to retrieval failures.
  • Methods like HyDE (Gao et al., 2022), Query2Doc (Wang et al., 2023), and QA-RAG (Kim and Min, 2024) try to address the alignment issue by leveraging LLMs to rewrite the user query into a form more aligned with the content space. These approaches implicitly approximate the retrieval function as $f^{*}(c,q_{0})\approx c^{t}p_{LLM}(q_{0})$, where $p_{LLM}(q_{0})$ represents the LLM’s output conditioned on the original query $q_{0}$.The effectiveness of these methods to indirectly align query and content space hinges on the LLM’s ability to generate representations that are both semantically rich and well-aligned with the content embeddings, which can be inconsistent and resource-intensive.
  • Adapter based methods (e.g. LlamaIndex222 https://docs.llamaindex.ai/en/stable/examples/finetuning/embeddings/finetune_embedding_adapter.html) approximates the retrieval function by $c^{t}U(q)$ where $U$ is an adapter function which must be learnt online, for example, with a single layer linear layer or even a deep neural net. However, acquiring sufficient training data for optimal adapter performance can be challenging. An alternative approach involves training separate, specialized embedders for queries and content (Karpukhin et al., 2020), aiming to learn inherently aligned representations.
  • Similar to adapter-based methods, this approach involves training or fine-tuning embedding models to generate representations specifically optimized for the retrieval task. Given a dataset of query-content pairs, these methods learn an embedding function $e(.)$ that maximizes the similarity between relevant queries and content. The retrieval function is then approximated as $f^{*}(c,q_{0})\approx e(c)^{t}e(q_{0})$. Commercially available tools, such as those on VertexAI [^3] https://cloud.google.com/vertex-ai/generative-ai/docs/models/tune-embeddings, offer pre-trained or fine-tuned embedding models for specific domains and tasks. This method essentially tries to handle alignment in a supervised way. However, this approach requires a substantial amount of labeled data for training, which can be costly and time-consuming to acquire, especially in specialized domains like healthcare.
We have intentionally excluded methods like re-ranking from this analysis. While re-ranking techniques can be applied subsequent to any of the aforementioned retrieval strategies, they don’t fundamentally alter the initial embedding-based retrieval process. Re-rankers typically leverage LLMs to evaluate the retrieval function $f^{*}(c,q_{0})$ on a smaller subset of documents $\mathcal{\tilde{C}}\subset\mathcal{C}$ that have been pre-selected using embedding similarity. This allows for a more precise approximation of $f^{*}(c,q_{0})$ but only within the confines of the initially retrieved subset. Determining the optimal subset for re-ranking remains an open challenge. Various frameworks, such as those offered by cohereAI [^4] https://cohere.com/rerank/, implement re-ranking mechanisms.
Finally, in the context of RAG – i.e. identifying relevant content for a downstream generation – we need only worry about finding some $c$ s.t. $f^{*}(c,q_{0})=1$. This simplifies the problem as we need to find contents that score high values following the chosen approximation function, rather than accurately estimating the entire function.
3.3 Misalignment
As many have previously noted (Gao et al., 2022; Wang et al., 2023; Kim and Min, 2024), directly using cosine similarity for retrieval within RAG systems can be problematic since these representations often reside in different semantic spaces, leading to misaligned retrieval. To clearly illustrate this, consider the following sentences and their corresponding embeddings:
  • $q_{1}=$ “Where was Queen Elizabeth born?”
  • $q_{2}=$ “Where was the Queen of Spain born?”
  • $c_{1}=$ “The former monarch of the British empire was born in London. Corgis are amazing fun dogs.”
  • $c_{2}=$ “The Spanish monarch was born in Madrid.”
  • $c_{3}=$ “The Queen of Spain was born in May.”
[…]
To overcome this alignment issue, some methods propose leveraging an LLM to draft a fake document or a candidate answer from the LLM’s implicit knowledge – see HyDE (Gao et al., 2022), Query2Doc (Wang et al., 2023), see QA-RAG (Kim and Min, 2024) – with a hope that these transformations and their embeddings would better align with those in $\mathcal{C}$ given we are now comparing documents to documents (or similarly answers to documents). These approaches are valuable as they recognize the misalignment, but are challenging to evaluate in an offline setting in a principled way. Their success depends not only on the LLM’s ability to comprehend the query and generate a relevant response (which may be challenging in specialized applications) but also on its capacity to produce answers that are semantically aligned with the specific content within the RAG system’s knowledge base.
We now connect this back with the ideal retrieval objective defined above. As we illustrated, when the inherent similarity between query and content embeddings fails to reflect their true relevance, the retrieval process is fundamentally hindered, often resulting in the retrieval of irrelevant content. This underscores the need for a more robust alignment strategy within RAG systems, one that can accurately capture the semantic relationship between queries and content, regardless of their inherent embedding representations. In the following section, we introduce QB-RAG, a novel approach that directly addresses this challenge by aligning incoming queries with a pre-computed set of questions derived from the content base and performing retrieval within a shared query space.
3.4 QB-RAG
To address the semantic misalignment inherent in traditional RAG retrieval, we propose Query-Based Retrieval Augmented Generation (QB-RAG). Our approach leverages a pre-computed set of questions, $\mathcal{Q}$, derived from the content base $\mathcal{C}$. These questions are generated offline as described in Section 2.2, mitigating concerns about online computational overhead. We not that efficient retrieval algorithms render the increased knowledge base size introduced by incorporating $\mathcal{Q}$ negligible. In contrast to methods relying on online LLM calls for query rewriting, QB-RAG shifts this computational burden offline. This distinction is significant, as online rewriting necessitates serial LLM invocation, potentially introducing latency detrimental to user experience. Furthermore, QB-RAG’s direct alignment within the query space offers a more transparent and interpretable retrieval process compared to the implicit alignment strategies of LLM-based rewriting techniques.
3.4.1 Vanilla QB-RAG
Our vanilla approach of QB-RAG first generates an extensive set of questions that are known to be answered by the content by initializing $\mathcal{Q}$ and $A$. This operation can happen offline and upon uploading new documents to our content base. Second, for an online query, QB-RAG searches for similar question (resp. questions) within $\mathcal{Q}$ by finding $\arg\min_{\mathcal{Q}}d(q_{0},q)$ (resp. $\arg\min_{\mathcal{S}\subset\mathcal{Q}}\sum_{q\in\mathcal{S}}{d(q,q_{0})}$). Given we are now comparing questions to questions, we expect the distance measure to be calibrated given the comparison is also aligned. Third upon retrieving similar questions, the associated contents are fetched, and fed to the generative LLM like other RAG systems (after dropping duplicate contents if necessary). We provide the full details in Algorithm 1.
Algorithm 1 Vanilla QB-RAG
$\mathcal{C},\mathcal{Q},A$ (or $A^{*},\hat{A}$), new query $q_{0}$, target number of retrievals $k\leq M$
Similarities $z\leftarrow Q^{t}q_{0}\in\mathbb{R}^{N}$
Sort $z$ by descending values and sort $\mathcal{Q}\&A$’s columns accordingly.
$\mathcal{S}\leftarrow\{\}$
for $j\in[1..N]$ do
     $i\leftarrow i$ s.t. $A_{ij}=1$ identify the content associated associated to the $j$-th question.
     Note, if $A$ is non-sparse, we break ties by favoring new content, then by number of associated questions, then at random.
     $\mathcal{S}\leftarrow\mathcal{S}\cup\{c_{i}\}$ if $c_{i}\notin\mathcal{S}$
     If $\lvert\mathcal{S}\rvert=k$ then break
end for
Return $\mathcal{S}$
3.4.2 Oracle and Approximate QB-RAG
We now turn our attention to the best possible retrieval system using a query based alignment and retriever. We previously noted that the matrix $A$ indicating which content generated which question, is actually a sparse partial observation of $A^{*}$ indicating which content is relevant to a question. For this section we will assume that we have access to the oracle matrix $A^{*}$ or some approximation of it denoted $\hat{A}$. For simplicity, we describe the algorithms with ${A}^{*}$, but is also applicable to $\hat{A}$.
Before getting into the algorithmic details, let us first motivate the use of this matrix $A^{*}$. We note that while computing the matrix completely would involve $N\times M$ LLM calls, these can all happen offline. Further depending on the dataset at hand, this may not even prohibitive in cost. E.g. in our dataset, this would lead to more than 5000 offline calls. Let us point out though that exactly computing $A^{*}$ is not necessary, and some approximation of it could also lead to improved retrieval. In fact, we propose two ways of computing the estimate $\hat{A}$:
  1. 1.
    Only compute matrix entries $A^{*}_{ij}$ such that $c_{i}^{t}q_{j}\geq\lambda$ where lambda is some threshold, which can be chosen in order to compute only some percentile of entries. This typically filters out combinations of contents and queries that are likely not relevant to each other.
  2. 2.
    After computing a set of entries (whether chosen at random or according to the above rule), we can rely on matrix completion techniques to infer and approximate the remainder of the matrix. Matrix completion is highly efficient on moderate to large matrix sizes.
Given the oracle $A^{*}$ or some estimate $\hat{A}$ thereof, we can adapt Algorithm 1 to incorporate the non-sparse nature of $A^{*}$ or $\hat{A}$. Specifically, since multiple content pieces may be associated with a single question, we introduce tie-breaking rules within the algorithm that first prioritize newer content (especially relevant to healthcare). We then prioritizes contents capable of answering a diverse range of questions within $\mathcal{Q}$. This reflects our intuition that a such content is likely to be more broadly relevant.
One could also explore incorporating the many-to-many relationship by leveraging a weighted majority vote into the retrieval step, rather than simple tie breaking rules. We defer this to future research as well.
4 Benchmark
4.1 Methods
This section outlines the retrieval methods evaluated in our benchmark. To ensure a fair comparison, each method retrieves the same number of documents, denoted by $k$, for a given query (the value of $k$ is varied across experiments). All LLM generations utilize Gemini-Pro and text embeddings are generated using textembedding-gecko – both are commercially available tools.
  1. 1.
    Naive RAG (Lewis et al., 2020). This baseline approach retrieves content based on the cosine similarity between the embedded query and candidate content using identical embedder, a standard practice in RAG systems Finally the query and retrieved contents are fed into an answer generator LLM and the LLM generates a response.
  2. 2.
    HyDE (Gao et al., 2022). HyDE addresses query-content misalignment by first generating a hypothetical document that answers the query. Specifically, an instruction tuned LLM is prompted to generate a hypothetical pseudo-document by setting the instruction “Write a paragraph that answers the question …” – this pseudo-document is not real. Second this pseudo-document and its embeddings are used for retrieval by searching for similar real contents, offloading the relevance modeling to the LLM’s generation.
    Query2Doc (Wang et al., 2023). Similarly to HyDE, Query2Doc employs few-shot prompting to generate a hypothetical pseudo-document that would answer the query. It concatenates the initial query with the synthetic document (generated similar to HyDE) and then performs dense retrieval using the embeddings of this concatenated text. Given the similarity with HyDE, we only implement HyDE in our benchmark.
  3. 3.
    QA-RAG (Kim and Min, 2024). In their methodology, they fine tune a foundation LLM offline using a golden set of Q&A. Upon receiving a new query online, they generate a pseudo-answer using the fine tuned LLM. They then retrieve contents using the query embeddings only, and other contents using the pseudo-answer’s embeddings only. A Reranker is then deployed to select only the most relevant contents among this dual set. In our implementation of QA-RAG, we use Gemini-Pro out of the box to generate a pseudo-answer, and retrieve $k$ documents using the pseudo-answer only, without a downstream reranker.
  4. 4.
    Vanilla QB-RAG: We run the simplest version of our method (see Algorithm 1) using the matrix $A$ which directly encodes “which content generated which question”. We believe that having access to the oracle matrix $A^{*}$ or some approximation of it would be even more promising. Furthermore, given this method is strongly dependent on the questions generated offline (in section 2), we parameterize QB-RAG-${\bar{m}}$ where $\bar{m}$ measures the average number of questions generated per content. Notably, $\bar{m}$ indicates the “coverage” extracted from each content, implying that higher coverage should lead to improved retrieval quality and, consequently, better-generated answers. We decrease $\bar{m}$ by down-sampling the set of questions in our experiments, to effectively reduce the coverage of our question base. The maximum value from our generation is $\bar{m}=8$ which includes our entire question set.
We point out that the methods presented above are not mutually exclusive and can be combined to potentially achieve improved performance, as illustrated by QA-RAG (Kim and Min, 2024). Specifically, we could retrieve contents using a combination of above methods and/or associated approached for generating embeddings, and subsequently leverage a re-ranker to select the most relevant documents. However, our experiments focus on isolating and evaluating the efficacy of each method in isolation.
Finally, we also highlight several other advancement introduced in RAG: 1. Modifications prior to retrieval by rewriting or breaking down the query (e.g. ITER-RETGEN (Shao et al., 2023)) in order to incorporate contextual information; 2. enhancements during retrieval by fine tuning the embedder (Karpukhin et al., 2020), training an adapter, retrieving diverse documents using MMR; 3. post retrieval refinements by reranking (or filtering, see (Nogueira et al., 2019)) the contents retrieved using an evaluator LLM, alternatively building a mixture LLM distributions (see REPLUG (Shi et al., 2023)), or even fine tuning the generator LLM to better utilize the retrieved contents (see RA-DIT (Lin et al., 2023)). While these techniques hold potential for enhancing retrieval and answer generation across all evaluated methods, we omit them from our analysis to maintain a focused comparison. This approach ensures an independent evaluation of the core retrieval mechanisms, as these enhancements can be incorporated without fundamentally altering the underlying methods.
4.2 Answer Generation
After retrieving relevant content using the methods described above, we employ an LLM (Gemini-Pro) for answer generation. For each question, the retrieved content is provided as context to the LLM along with the following prompt to generate the answer:
1Use only the provided pieces of context to answer the question at the end. Think step-by-step and then answer. Respond in 3 to 6 short sentences.
2
3Do not try to make up an answer:
4- If the context does not contain enough relevant information to determine an answer to the query, say "I cannot determine the answer to that."
5- If the context is empty, just say "I do not know the answer to that."
6
7Answer in an empathetic and positive tone. Do not use phrases such as "According to the context", and directly answer the question.
8
9Contexts: {contexts}
10Question: {question}
11Answer:
Listing 3: Base prompt for answer generation
4.3 Test Data Generation
We construct two distinct test sets designed to assess the performance of the various retrieval methods under different conditions. Each test set consists of questions answerable by our content base, ensuring relevance to the task. Crucially, no test question appears verbatim within the pre-generated question set used for retrieval. All test questions are generated using Gemini-Pro..
  1. 1.
    Rephrase: To generate this test set, we prompted an LLM to rephrase each of the 4.8k questions in the knowledge base. Then, 500 of those were randomly sampled to yield the first test set. This approach was done to simulate scenarios where the knowledge base contains a comprehensive set of questions. In such scenarios, the intents of new incoming questions are more likely to be represented in the knowledge base.
  2. 2.
    Out-of-Distribution: To generate the second test set, we prompted an LLM to generate a new question for each of the 630 contents. The LLM was instructed to not generate a question that already exists in the question knowledge base for that content. Then, the 630 newly generated questions were filtered via the answerability model from Section 2 to ensure the new questions were indeed answerable by the corresponding content. After filtering, 305 newly generated questions made up the second test set.
4.4 Metrics
To rigorously assess the performance of our proposed QB-RAG method, we employ two distinct sets of metrics: those evaluating the quality of content retrieval and those assessing the quality of the generated answers. While QB-RAG’s primary objective is to enhance retrieval, we hypothesize that this improvement will translate to better-generated answers. Given the emphasis on reliability and safety in healthcare applications, we carefully selected answer quality metrics tailored to this domain. Since our content base represents a vetted source of trusted information, we prioritize answer faithfulness, penalizing any unwarranted extrapolation or information not grounded in the provided content. Additionally, we recognize the critical importance of a model’s ability to decline answering when the content lacks clarity, a crucial aspect of mitigating potential risks in healthcare settings.
4.4.1 Retrieval Evaluation
  • Exact Recovery Rate: This metric measures the percentage of test questions for which the retrieved set of $k$ documents includes the exact content piece used to generate the original question. We acknowledge that while exact recovery represents retrieval of “golden content”, multiple contents might offer relevant information for a single question. Therefore, we incorporate additional relevance measures described below.
  • Auto-evaluator Relevancy Rate: This metric addresses the limitations of relying solely on exact matches by leveraging the answerability model (Section 2) to gauge content relevance. Specifically, it calculates the percentage of test questions for which at least one retrieved document is deemed relevant by the answerability model. This automated assessment has demonstrated strong correlation with human judgments as as described in section 2.
  • Average Re-ranker Relevancy Score: To evaluate the relevance of each retrieved document individually, we employ the BGE re-ranker (Xiao et al., 2023). This method assigns a relevance score (higher indicating greater relevance) between each retrieved document and its corresponding test question. For each test question, we take the maximum relevance score among the $k$ retrieved documents and average these maximum scores across all test questions. The effectiveness of the BGE re-ranker in relevance scoring has been previously established in RAG research (Kim and Min, 2024).
4.4.2 Answer Evaluation
  • Answer Guideline Adherence Rate: This metric measures how well a generated answer aligns with a predefined guideline derived from a “golden answer.” The assessment involves a three-step process. Initially, an LLM generates a “golden answer” for each test question using the associated content. Subsequently, another LLM analyzes this golden answer to create a guideline outlining the key elements an accurate response should contain. Finally, a third LLM evaluates the candidate answer against this extracted guideline, assigning a score from 0 to 1 based on the extent to which the answer covers the guideline’s key points. While this guideline-based approach aims to capture the nuances of answer quality, it relies heavily on the accuracy of both the golden answer and the extracted rubric. To provide a more robust and multifaceted evaluation, we introduce additional metrics that directly assess distinct aspects of answer quality below.
  • Answer Relevancy Rate: This metric evaluates whether a generated answer directly address the user’s question and provides a self-contained response. An LLM classifies each answer as either relevant (YES) or not relevant (NO) to the corresponding test question, focusing solely on the relevance, not the accuracy or factual correctness of the answer.
  • Answer Faithfulness Rate: This evaluates weather the content supplied to the LLM during generation supports the generated answer. An answer is deemed faithful if any portion of the provided content supports it, regardless of the presence of irrelevant content. This component is crucial to ensure that the answer generated is grounded on our content, a crucial feature for many healthcare applications. Ensuring the answer is grounded on the supplied content will mitigate the hallucination risk. Note that declined answers are considered not faithful in our definition.
  • Answer Declined Rate: This metric assesses whether the LLM declined to answer the question. This may occur because our answer generation couldn’t find relevant information in the content (we specifically prompt the LLM to only answer questions when there is a retrieved content.)
5 Experimental Results
In this section we discuss the performance of our methods, QB-RAG-8 (resp. QB-RAG-2) where the knowledge base has an average of 8 (resp. 2) queries per content, which are compared against the incoming query for retrieval. We assess the impact of QB-RAG on both retrieval efficacy and the quality of generated answers.
5.1 Effect of QB-RAG on Retrieval Efficacy
On the Rephrase test set, where the knowledge base is expected to contain questions semantically similar to the test questions, QB-RAG-8 consistently outperforms all benchmark methods. As shown in table 2, QB-RAG-8 nearly doubles the exact recovery rate compared to other methods (e.g., from 45% to 89% when retrieving a single document). This suggests that a comprehensive, query-aligned knowledge base, as constructed by QB-RAG-8, substantially improves the retrieval of the exact source document. These impressive gains highlight the scenario where our knowledge base is comprehensive and covers quite broadly the extent of questions our documents can answer.
The exact recovery rate does not present the full picture. When analyzing the Auto-evaluator Relevancy and the Re-ranker score (refer section 4.4.1), our method still dominates the baselines when retrieving a single content. As measured by the LLM, the content we retrieve are relevant 68% of the time, whereas traditional methods retrieve relevant content around 40% of the time. In table 2, we also include the average re-ranker relevancy score. QB-RAG-8 yields the highest average relevance score using the BGE re-ranker.
Increasing the number of retrieved documents to 3 further illustrates the efficacy of QB-RAG-8. The exact recovery rate reaches a near-optimal 97%, substantially higher than the baseline approaches. QB-RAG-8 also achieves almost 20% higher relevancy rate compared to the baselines. Similarly, the average re-ranker relevancy score is significantly higher for QB-RAG. These findings underscore the robustness of QB-RAG-8 in retrieving both the target document and a set of relevant documents.
Retrieval assessment Rephrase Exact Recovery Rate Auto-evaluator Relevancy Rate Avg. Re-ranker Relevancy Score
Number of retrieved docs 1 3 1 3 1 3
QB-RAG-8 0.89 0.97 0.68 0.76 1.21 1.65
QB-RAG-2 0.59 0.75 0.53 0.66 –0.25 0.92
Naive RAG 0.44 0.61 0.39 0.56 -0.96 0.35
QA-RAG 0.45 0.60 0.41 0.56 -1.52 -0.38
HyDE 0.47 0.63 0.41 0.56 -1.14 -0.09
Table 2: Retrieval performance (higher is better) of methods on Rephrase, where documents were retrieved given rephrased questions from the content base.
On the more challenging Out-of-Distribution test set, QB-RAG-8 maintains its advantage, albeit with smaller gains. Table 3 shows that QB-RAG-8 improves the exact recovery rate by 1.3% to 6.2% compared to benchmark methods. Despite the adversarial nature of this test set, where incoming questions are intentionally dissimilar to the training set, QB-RAG-8 consistently retrieves the correct source document more often.
Furthermore, QB-RAG-8 demonstrates a significant improvement in relevancy. The auto-evaluator relevancy rate shows gains of 8% to 15% over baseline methods, indicating that QB-RAG-8 effectively identifies relevant content even when the query distribution shifts. These improvements in retrieval quality directly benefit downstream answer generation, as more relevant content is likely to lead to more accurate and informative answers. This is further supported by the consistent gains observed in the average relevance scores from the BGE re-ranker.
Retrieval assessment Out-of-Dist Exact Recovery Rate Auto-evaluator Relevancy Rate Avg. Re-ranker Relevancy Score
Number of retrieved docs 1 3 1 3 1 3
QB-RAG-8 0.53 0.72 0.58 0.75 0.47 1.68
QB-RAG-2 0.42 0.57 0.50 0.64 -0.54 0.78
Naive RAG 0.50 0.68 0.50 0.64 -0.03 1.25
QA-RAG 0.51 0.66 0.47 0.60 -0.48 0.53
HyDE 0.52 0.70 0.48 0.65 -0.01 1.20
Table 3: Retrieval performance (higher is better) of methods on Out-of-Distribution, where documents were retrieved given newly generated questions.
5.2 Effect of QB-RAG on Generated Answer Quality
We now examine how the improved retrieval accuracy of QB-RAG translates to the quality of generated answers. Recall that the Rephrase test set favors our query-based approach as test questions are semantically similar to those in our generated knowledge base.
As shown in Table 4, both QB-RAG-8 and QB-RAG-2 consistently outperform the benchmark methods on all answer quality metrics for the Rephrase test set. Notably, QB-RAG-8 achieves an 84% answer faithfulness rate, significantly surpassing the 62%-68% rates of the baseline methods. This suggests that by accurately retrieving the most relevant content, QB-RAG enables the LLM to generate answers that are well-grounded in the provided information. Additionally, QB-RAG achieves the highest accuracy rate, indicating its answers effectively address the key elements outlined in the pre-defined guidelines (refer to Section 4.4.2).
Answer assessment Rephrase Declined Rate Faithfulness Rate Relevancy Rate Accuracy Rate
QB-RAG-8 .12 .84 .83 .79
QB-RAG-2 .21 .74 .73 .73
Naive RAG .30 .67 .66 .68
QA-RAG .33 .62 .62 .62
HyDE .30 .68 .66 .67
Table 4: Answer quality of methods on Rephrase, where documents were retrieved given rephrased questions from the content base. 3 documents retrieved.
These patterns hold on the more challenging Out-of-Distribution test set, as seen in table 5. There QB-RAG-8 achieves 78% faithfulness, fairing higher than the benchmarks 68%-74%. However QB-RAG-2 under-performs, highlighting its inability to generalize to new questions.
Answer assessment Out-of-Dist Declined Rate Faithfulness Rate Relevancy Rate Accuracy Rate
QB-RAG-8 .17 .78 .77 .69
QB-RAG-2 .29 .66 .65 .63
Naive RAG .27 .72 .70 .64
QA-RAG .29 .68 .67 .63
HyDE .24 .74 .72 .68
Table 5: Answer quality of methods on Out-of-Distribution, where documents were retrieved given newly generated questions. 3 documents retrieved.
An answer can be deemed unfaithful either because the answer is not grounded, or because the LLM declined to answer. We note that the higher faithfulness rate of QB-RAG-8 is largely due to improved retrieval, resulting in a decline rate of only 12% on the Rephrase test set and 17% on the Out-of-Distribution test set, compared to significantly higher rates for the baselines. We observe a slight increase in answers that are not grounded, from $\approx 2\%-3\%$ on the Rephrase test set to $\approx 3\%-5\%$ on the Out-of-Distribution test set for QB-RAG. Importantly, we achieve the lowest unfaithfulness rate (3%) on the Rephrase test set with QB-RAG-8, further underscoring the value of generating a comprehensive question set for optimal coverage.
While QB-RAG significantly improves answer faithfulness through better retrieval, it’s worth noting that the groundedness aspect of faithfulness rate is ultimately determined by the capabilities of the answer generation module itself. Further enhancements to answer faithfulness could involve techniques like preference modeling and RLHF where the LLM is specifically for this objective (Bai et al., 2022).
5.3 Effect of Coverage of Generated Questions
To examine the relationship between the breadth of the generated question set and QB-RAG’s effectiveness, we compare the performance of QB-RAG-2 and QB-RAG-8. This analysis reveals a strong dependence on question coverage.
On the Rephrase test set, QB-RAG-2, despite its reduced question set, still surpasses other retrieval methods (table 2). However, the magnitude of improvement is noticeably smaller compared to QB-RAG-8. For instance, QB-RAG-2 shows a 10-15% improvement in exact recovery rate over baselines, whereas QB-RAG-8 nearly doubles the exact recovery rate. This pattern also holds for relevancy rate and average re-ranker scores, indicating that a larger, more comprehensive question set translates to more effective retrieval.
The Out-of-Distribution test set (table 3) reveals a more pronounced impact of question base coverage. Here, QB-RAG-2’s performance drops below that of some benchmark methods, with the exact recovery rate decreasing by 7-12%. Interestingly, QB-RAG-2 still achieves comparable performance on relevancy-based metrics, suggesting that even a limited question base can partially capture relevant content, but may not pinpoint the exact source document as effectively.
This sensitivity analysis clearly shows that the performance of QB-RAG methods are very much tied to our ability to generate a comprehensive set of questions from RAG knowledge base. As we increase the number of generated questions in our database, we expect to see better retrieval, higher exact matches and higher relevance. Beyond simply increasing the number of questions, improving their diversity may be another way for further improving QB-RAG’s performance. By generating a more diverse set of questions for each content piece, we can capture a broader range of semantic nuances and user intents. This is particularly relevant given that modern retrieval algorithms efficiently handle large knowledge bases, making quantity less of a limiting factor in doing efficient retrieval.
This sensitivity to question coverage is further evident in the quality of the generated answers (Tables 4 and 5). QB-RAG-8 consistently leads to more faithful, relevant, and accurate answers compared to QB-RAG-2, directly reflecting the differences observed in their retrieval performance. These findings highlight that QB-RAG’s success in downstream tasks is fundamentally linked to its ability to construct and leverage a comprehensive and diverse question set that effectively captures the content and semantic nuances of incoming queries within the RAG knowledge base.
5.4 Case Study
To more concretely demonstrate the advantages of QB-RAG in our application, we present an example query from the out-of-distribution test case. The results of Naive RAG, QA-RAG and HyDE for the given test query are as follows.
Test query:
What is the difference between soluble and insoluble fiber and how do they affect the body?
Retrieved content from all non QB-RAG methods:
Fiber is the “roughage” in plants that gives them their shape. Your body can’t digest it (it just passes through), so why do we need to eat it? You’ve probably heard that fiber helps keep you regular. Fiber can also help lower your cholesterol, manage your blood glucose, and keep you feeling full longer. And it can help you lose weight. How much fiber is enough? For women, experts recommend 25 grams of fiber each day. Men need about 35 grams. Ready to add more fiber to your diet? Try a couple of these tips this week: Choose 100% whole-grain bread or tortillas instead of white bread or white-flour tortillas. Eat brown rice instead of white rice. Fill half your plate with low-carb vegetables. Grab an apple instead of applesauce. Swap in black bean or navy bean soup instead of chicken noodle soup. Choose oatmeal or oat bran instead of corn flakes or puffed rice cereal. Munch on popcorn or kale chips instead of pretzels. Consider trying a fiber-enriched food. A fiber-enriched breakfast bar has around one-third of your daily fiber! A quick note: Fiber is wonderful, but add it to your diet slowly. Your body may need a little time to adjust. As you add fiber, drink more fluids to keep it moving through. If you have any problems with gas or bloating, there are over-the-counter that can help. Which sources of fiber look good to you?
Answer generated using retrieved content from all non QB-RAG methods:
I cannot determine the answer to that. The provided context does not mention the difference between soluble and insoluble fiber or how they affect the body.
Given the query about soluble and insoluble fiber, Naive RAG, QA-RAG and HyDE all retrieved a content about fiber as the top content. While on topic, the content does not contain information to differentiate the effect of soluble and insoluble fiber on the human body, leading to the answer generator declining to answer. The results of QB-RAG for the same test query are as follows.
Retrieved query by QB-RAG:
How does soluble fiber help the body?
Mapped content of retrieved query by QB-RAG:
Steel-cut oatmeal (also called Irish oatmeal) is an easy, healthy, high-fiber food. Steel-cut oats have more fiber than rolled or instant oats. Steel-cut oats have both soluble and insoluble fiber. Soluble fiber helps lower cholesterol and keeps blood glucose from rising quickly. Insoluble fiber helps move food through your body, which can keep your bowels regular. The fiber in oatmeal also makes it a very filling food. You can eat a small portion size and feel full longer, which is great for weight loss. Want to boost your oatmeal flavor? Try these ideas: Cinnamon Slivered almonds or other nuts (2 Tbsp) Pumpkin pie spices without sugar Almond butter (1 Tbsp) or other nut butter Powdered peanut butter Powdered protein powder (mix in after cooking for a protein and flavor boost) Stevia (sugar-free sweetener) A dash of vanilla extract Tips: Oatmeal has carbs so keep a small portion size in mind. Milk adds carbs. Consider using a cup of regular milk, unsweetened almond milk, or unsweetened coconut milk as low-carb options. Try eating oatmeal with a protein food (like eggs, nuts, or peanut butter) for a balanced meal.
Answer generated using mapped content of retrieved query by QB-RAG:
Soluble fiber helps lower cholesterol and keeps blood glucose from rising quickly, while insoluble fiber helps move food through your body, which can keep your bowels regular.
Given the same test query, QB-RAG first retrieved a pre-generated query about soluble fiber stored in the knowledge base. It then maps the retrieved query to the content that was used to generate the query. Given the content, the answer generator was able to provide a response to the test query. Note that while the retrieved content is largely about steel-cut oatmeal, it contains information about soluble and insoluble fiber and their effect on the body. On the surface, it may seem that the content retrieved by the other methods is more on topic, QB-RAG was able to pinpoint the exact content that contains the needed specific information with the help of the pre-generated query.
6 Discussion
This paper introduces QB-RAG, a novel approach for enhancing the retrieval phase of RAG systems. QB-RAG addresses the inherent semantic misalignment between user queries and knowledge bases by pre-generating a comprehensive set of potential questions directly from the content. We then leverage efficient vector search to map incoming online queries to these pre-computed questions, facilitating a more accurate and aligned retrieval process. While QB-RAG requires a significant offline computational investment for question generation, this trade-off is strategically advantageous. Unlike conventional methods that rely on online LLM calls for query rewriting or enhancement, QB-RAG shifts this computational burden offline. This distinction is crucial for minimizing latency in real-time applications, as online LLM invocations can significantly impact the user experience. Furthermore, QB-RAG benefits from the efficiency of highly optimized retrieval algorithms, even when handling large question sets, ensuring rapid content retrieval for online queries. We have provided a theoretical motivations for QB-RAG and we have also suggested potential avenues for further improving QB-RAG, such as expanding the question set for greater coverage or leveraging matrix completion techniques to infer a complete answerability matrix.
To assess QB-RAG’s effectiveness, we designed two distinct test sets: the Rephrase set, where test questions are semantically similar to those in our pre-generated question base, and the Out-of-Distribution set, where questions are intentionally crafted to be dissimilar. These test sets represent cases designed to evaluate QB-RAG’s robustness under varying conditions, including a potential distribution shift. Our empirical results demonstrate that QB-RAG consistently outperforms benchmark methods across both test sets. While QB-RAG significantly improves exact recovery and relevance of retrieved documents, its impact extends beyond these initial retrieval metrics. Downstream answer generation using QB-RAG-retrieved content scores higher on key quality metrics, including faithfulness, relevancy, and adherence to pre-defined guidelines. Crucially, answers generated using QB-RAG are more frequently grounded on our trusted source of content, which is key to deliver reliable, up to date information to patient queries.
Our sensitivity analysis shows that effectiveness of our approach is highly tied to coverage of extracted questions. Beyond increasing the raw number of questions, we also suggest favoring a diverse set of questions during the generation phase. This can be done potentially with prompt engineering, supervised fine-tuning or RLHF frameworks. Metrics like BERTScore (Zhang et al., 2019) and Word Mover’s Distance (Kusner et al., 2015) can be used to evaluate the diversity of generated questions from the RAG knowledge base. Finally, an LLM could also be potentially used as an auto-evaluator for measuring diversity of generated questions. We note that retrieval frameworks are highly optimized even for large question bases, thus, we don’t expect quantity of generated questions as a bottleneck for many real-world applications.
Despite its novelty and effectiveness, our work has several limitations that require further study. First, our evaluation focused on a curated healthcare content base consisting of well-crafted, concise pieces. Assessing QB-RAG’s performance on larger, more diverse datasets, including those from other domains and containing less structured or longer-form content, remains an important area for future research.
Another limitation stems from our reliance on LLM-based evaluation metrics in the current study. While LLMs have shown promise in assessing answer quality according to predefined guidelines, the real-world deployment of QB-RAG in healthcare settings necessitates thorough expert review. Direct patient interactions demand the highest level of scrutiny to ensure the accuracy, safety, and reliability of generated information. Our automated evaluation serves as a valuable first step, but expert validation remains paramount.
Finally, it’s crucial to acknowledge a broader challenge inherent to all RAG systems, particularly in fast-changing domains like healthcare: the quality of any RAG system’s output is fundamentally dependent on the quality and recency of its underlying knowledge base. In healthcare, where medical consensus and best practices are constantly evolving, content needs to be regularly reviewed and updated to reflect the latest advancements. This underscores the importance of developing robust mechanisms for ongoing knowledge base maintenance and ensuring alignment with current clinical guidelines, regardless of the specific retrieval methods employed.
Despite these limitations, our work demonstrates the potential of QB-RAG for enhancing retrieval accuracy and answer groundedness in RAG systems, paving the way for more reliable and trustworthy LLM applications in healthcare. Beyond direct patient interaction, QB-RAG could also be valuable for augmenting coaching and provider resources within digital health platforms. For instance, QB-RAG could be used to generate draft answers that healthcare professionals can then review and refine, ensuring both efficiency and accuracy. Ultimately, real-world deployment of QB-RAG should be carefully considered, taking into account specific use cases, potential risks, and the need for appropriate expert oversight.